OpenStudy (anonymous):

A sample of 100 wood and 100 graphite tennis rackets are taken from the warehouse. If 6 wood and 14 graphite are defective and one racket is randomly selected from the sample, find the probability that the racket is wood or defective.

OpenStudy (anonymous):

total 200 rackets........ P(A U B) = P(A) + P(B) - P(AnB) 100/200 + 20/200 - 6/200 = 114/200